OpenStudy (anonymous):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 9 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 15 over 20 11 over 20 6 over 20 5 over 20

OpenStudy (anonymous):

i think A

OpenStudy (anonymous):

hey do u still need help??

Mehek (mehek14):

\(\dfrac{15}{20}\) is he probability of getting red and blue

Mehek (mehek14):

to find the probability of green, subtract \(\dfrac{20}{20}-\dfrac{15}{20}\)

Mehek (mehek14):

so what's 20 - 15?

OpenStudy (anonymous):

5

Mehek (mehek14):

so the probability of getting green is \(\dfrac{5}{20}\)

OpenStudy (anonymous):

Ohhhhhhhhh i get it now

OpenStudy (anonymous):

The answer is D!!!! :D